# Hydrogen liquid-liquid transition from first principles and machine learning

### v1

@@ Conclusions.
The physical picture provided by the MACE model qualitatively agrees with Ref. [34] on the supercritical nature of the LLT, even though their model could not accurately reproduce the crystallization regime, and predicted a Widom line (estimated from the maximum of the isobaric specif heat $c_{p}$ ) far from the AIMD results (see Fig. 4). The first-order character of the transition observed in other studies may be explained by the large fluctuations expected close to the Widom line, which could lead to the incorrect identification of density jumps and pressure plateaux in the EOS for too short simulation times. Additional path integral MD simulations carried out with the MACE model indicate that the conclusions about the crossover nature of the LLT are not changed by the inclusion of nuclear quantum effects, as reported in the SM [49].
@@
A version of the MACE code implementing the modified loss function employed in this work can be found at [63]. The model, training set, and simulation results are available at [64].
@@ Acknowledgments.
K.N. acknowledges financial support from MEXT Leading Initiative for Excellent Young Researchers (Grant No. JPMXS0320220025). M. C. thanks the European High Performance Computing Joint Undertaking (JU) for the partial support through the ”EU-Japan Alliance in HPC” HANAMI project (Hpc AlliaNce for Applications and supercoMputing Innovation: the Europe - Japan collaboration). This work has received funding from the European Center of Excellence in Exascale Computing TREX (Targeting Real chemical accuracy at the Exascale, grant 952165). The authors acknowledge discussions with Prof. David Ceperley, Prof. Michele Ceriotti, Prof. Gábor Csányi, Prof. Carlo Pierleoni, Dr. Thomas Bischoff, Dr. Marco Cherubini, and Dr. Abhishek Raghav. M.C. acknowledges GENCI for providing computational resources on the CEA-TGCC Irene supercomputer cluster under project number A0170906493 and the TGCC special session.

### v2

@@ Conclusions.
The physical picture provided by the MACE model qualitatively agrees with Ref. [34] on the supercritical nature of the LLT, even though their model could not accurately reproduce the crystallization regime, and predicted a Widom line (estimated from the maximum of the isobaric specific heat $c_{p}$) far from the AIMD results (see Fig. 4). The first-order character of the transition observed in other studies may be explained by the large fluctuations expected close to the Widom line, which could lead to the incorrect identification of density jumps and pressure plateaux in the EOS for too short simulation times. Additional path integral MD simulations carried out with the MACE model indicate that the conclusions about the crossover nature of the LLT are not changed by the inclusion of nuclear quantum effects, as reported in the SM [49].
@@
A version of the MACE code implementing the modified loss function employed in this work can be found at Ref. [63]. The model, training set, and simulation results are available at Ref. [64].
@@ Acknowledgments.
K.N. acknowledges financial support from MEXT Leading Initiative for Excellent Young Researchers (Grant No. JPMXS0320220025) and from by JST BOOST (Grant No. JPMJBY24F3). B.J. acknowledges support by the state of Baden-Württemberg through bwHPC. M. C. thanks the European High Performance Computing Joint Undertaking (JU) for the partial support through the ”EU-Japan Alliance in HPC” HANAMI project (Hpc AlliaNce for Applications and supercoMputing Innovation: the Europe - Japan collaboration). This work has received funding from the European Center of Excellence in Exascale Computing TREX (Targeting Real chemical accuracy at the Exascale, grant 952165). The authors acknowledge discussions with Prof. David Ceperley, Prof. Michele Ceriotti, Prof. Gábor Csányi, Prof. Carlo Pierleoni, Dr. Thomas Bischoff, Dr. Marco Cherubini, and Dr. Abhishek Raghav. M.C. acknowledges GENCI for providing computational resources on the CEA-TGCC Irene and IDRIS Jean-Zay supercomputer clusters under project number A0170906493 and the TGCC special session.